OpenStudy (anonymous):

Geometry :D Suppose f and g are two isometries such that f(A) = g(A), f(B) = g(B), and f(C) = g(C) for some nondegenerate triangle ΔABC. Show that f = g

OpenStudy (anonymous):

I feel like this should be easy, but still getting used to all the geometry theorems and what not.

OpenStudy (anonymous):

@freckles know anything about this stuff?

OpenStudy (anonymous):

@ganeshie8 Know any of this stuff?

ganeshie8 (ganeshie8):

idk much but i found this #3 http://www.ms.uky.edu/~droyster/courses/spring04/homework/restricted/Homework2%20Solns.pdf @ikram002p did non eucledean geometries a couple of years ago

OpenStudy (anonymous):

Ah, nice find! I found something else online, but it used a theorem that I couldnt use based on the sequence of my text. But alright, if ikram knows some of this stuff, Ill see what he can do when I have questions. I think I can just use what that link gives. Thanks :)

ganeshie8 (ganeshie8):

good luck! :)

OpenStudy (ikram002p):

hey first u need to know what isometries functions means, it means a rigid transforms function which maps a shape to some where else but without changing size in other way both origin and transforms are congruent.

OpenStudy (ikram002p):

lets do it in geometry style 

OpenStudy (ikram002p):

so since both are isometries we have 3 congruent triangles, now here is a thing why it called rigid transformation it since u do not change it size and shape , like triangle do not become a circle for example or dont become another triangle with same size but different sides that cannot happen.

OpenStudy (ikram002p):

now step two, it says if f(A) = g(A), f(B) = g(B), and f(C) = g(C), then show Show that f = g in euclidean geometry its ok if u wanna only graph it that works like this it would be a proof by itself 

OpenStudy (ikram002p):

now since you wanna it theoretically f(A)=g(A) g^-1o f(A)=g^_1 o g(A) =A so if u apply the function g^-1 o f on the other two points u got g^-1 o f(B)=B g^-1 o f(C)=C according to a theorem (idk if ts given in ur book or u need to prove it let me know in both cases ) g^-1o f must be the identity which means :- g^-1 o f(P)=P , for any P on the triangle g o g^-1 o f(P)= g(P) f(P)=g(P) which is done :D let me know if something is not clear enough, good luck !!